# Adapting Pre-Trained Vision Models for Novel Instance Detection and Segmentation

### v2

@@ Abstract
Novel Instance Detection and Segmentation (NIDS) aims at detecting and segmenting novel object instances given a few examples of each instance. We propose a unified, simple yet effective framework (NIDS-Net) comprising object proposal generation, embedding creation for both instance templates and proposal regions, and embedding matching for instance label assignment. Leveraging recent advancements in large vision methods, we utilize Grounding DINO and Segment Anything Model (SAM) to obtain object proposals with accurate bounding boxes and masks. Central to our approach is the generation of high-quality instance embeddings. We utilize foreground feature averages of patch embeddings from the DINOv2 ViT backbone, followed by refinement through a weight adapter mechanism that we introduce. We show experimentally that our weight adapter can adjust the embeddings locally within their feature space and effectively limit overfitting in the few-shot setting. This methodology enables a straightforward matching strategy, resulting in significant performance gains. Our framework surpasses current state-of-the-art methods, demonstrating notable improvements of 22.3, 46.2, 10.3, and 24.0 in average precision (AP) across four detection datasets. In instance segmentation tasks on seven core datasets of the BOP challenge, our method is around 4.5 times faster than the leading published RGB method and surpasses it by 3.6 AP. NIDS-Net is about 5.7 times faster than the top RGB-D method while maintaining competitive performance [^1] Project Page: https://irvlutd.github.io/NIDSNet/.
1 Introduction
@@
Novel Instance Detection and Segmentation (NIDS) is a crucial task in computer vision aimed at identifying and locating unseen instances in images or videos, given a few examples of each instance. It also has wide applications in robotics. For example, suppose a robot needs to grasp a specific, novel object instance from a cluttered desk given only a small number of multi-view template images of the object. Ideally, detectors for this task should provide precise bounding boxes and masks of the target instances given a query image. However, standard detection methods trained on specific classes often struggle to recognize new objects.
The current paradigm for NIDS typically encompasses the following steps: (1) generating proposals from a query image, (2) obtaining embeddings of the proposals and the object templates, and (3) matching the embeddings of proposals with those of the templates for identification. Recent work [49, 81, 66, 53] has utilized various open-world detectors to obtain object proposals. However, the use of models such as the Segment Anything Model (SAM) [47] or FastSAM [103] often results in the generation of region-based proposals rather than actual object proposals. This misidentification of regions as objects can lead to object identification issues. For example, a single object may be divided into multiple proposals or some background regions may be mistaken for foreground objects. These false alarms not only slow down the whole detection progress but also spoil the final results.
To embed proposals and templates, some existing works, e.g., [81, 66, 53], adopt the $cls$ token of DINOv2  [68, 20] while others, e.g., Li et al. [49], employ a 3D voxel representation. Ideally, for each specific unseen instance, the template embeddings from different views should be similar to each other but markedly different from the embeddings of other instances. Despite these efforts, the embeddings of different instances may remain similar, and so the challenges persist.
@@
To address the limitations of the existing methods, we propose a framework for NIDS, as depicted in Fig. 1. Initially, we utilize Grounding DINO [56] with the text prompt “objects” on a cluttered query image to obtain bounding boxes for foreground objects. This method capitalizes on the inherent objectness of Grounding DINO using specialized prompts rather than standard labels. Subsequently, we employ SAM to create masks within these bounding boxes, generating precise object proposals comprising both bounding boxes and masks. For instance embeddings, we first extract average foreground features [48] from the patch embeddings of the DINOv2 ViT backbone. We then apply adapters to enhance instance embeddings by clustering similar instances and distancing different ones.
@@
The adapters are trained using the template images only, where InfoNCE loss [67, 15] is applied to the feature embeddings after the adapters. To refine these embeddings, the CLIP-Adapter [29] introduces a residual vector added to the original embedding, which risks overfitting with a few training examples. This addition can spoil and destabilize the embeddings of background objects, disrupting the entire feature space. To mitigate this issue, we, instead, introduce a weight adapter that modifies the original embeddings by applying learned weights. Finally, we employ a straightforward matching approach [62], such as stable matching or using $argmax$, to assign instance labels to these proposals.
@@
Our approach has been validated on four detection datasets and seven segmentation datasets for NIDS. The framework significantly surpasses existing state-of-the-art methods: it demonstrates substantial increases in average precision (AP), with gains of 22.3, 46.2, 10.3, and 24.0 across these detection datasets. Moreover, in instance segmentation tasks on the seven core datasets of the BOP challenge [84], our method outperforms the best published RGB and RGB-D methods.
@@
  • We propose NIDS-Net, a unified framework for novel instance detection and segmentation that surpasses existing published methods across four detection datasets and seven segmentation datasets of BOP challenge.
@@
  • We leverage the objectness of the open-world detector Grounding DINO and region recognition of Segmentation Anything Model (SAM) to obtain object proposals.
@@
  • We introduce the Weight Adapter, a tool designed to refine embeddings within their feature space while discouraging overfitting. This approach yields increasingly significant improvements as the adapting of the feature space becomes more robust and stable.
2 Related Work
@@
Pretrained Models. Large-scale pretraining has demonstrated utility across various downstream tasks. Works such as [32, 91, 59, 24, 58, 10, 68] emphasize large-scale image representation learning, while [74, 50] focus on vision-language pretraining. Additionally, [3, 4, 96] address depth prediction, and [72, 73, 6, 87] tackle large-scale language modeling. Object detection is addressed by [79, 77, 43, 44, 31, 56], while image segmentation is the focus of [33, 89, 46]. Works such as [94, 52, 14, 76] explore text-to-3D generation, and [82, 1] address text-to-video synthesis. Finally, 6DoF pose estimation and tracking are tackled by [90, 21, 51, 88]. DINOv2 [68, 20] offers robust features to represent unseen instances. These foundational models offer diverse capabilities, with the challenge lying in effectively leveraging their wealth of knowledge for specific use cases, such as novel instance detection. In our work, we leverage pre-trained vision models for NIDS.
Open-World Object Detection (OWOD). OWOD tackles real-world scenarios where object detection algorithms are tasked not only with identifying known and labeled objects but also with effectively handling novel and unknown objects during inference. This sets OWOD apart from traditional object detection, which confines itself to detecting only predetermined object classes. A key challenge in OWOD lies in detecting and classifying unknown objects that were not part of the training data.
Recent approaches such as OWL-ViT [65], DITO [43], RO-ViT [44], YOLO-World [18], Sambor [31], and Grounding DINO [56] offer diverse solutions to address OWOD challenges. For example, OWL-ViT leverages a standard Vision Transformer [24] architecture with minimal modifications, incorporating contrastive image-text pre-training and end-to-end detection fine-tuning. DITO introduces an open-vocabulary detection approach based on detection-oriented image-text pretraining, replacing the classification architecture with a detector architecture during pretraining to better serve region-level recognition needs. Grounding DINO [56] utilizes features from DINO [10] to ground text queries. We employ Grounding DINO due to its strong performance. Similar to our work, Fang et al. [28] obtain object bounding boxes using the text prompt “a rigid object” with Grounding DINO.
Few-Shot Object Detection (FSOD). FSOD involves detecting occurrences of a novel class, unseen during training, given only a few template examples. For example, De-ViT [102] introduces a region-propagation mechanism for localization, transforming propagated region masks into bounding boxes using a learnable spatial integral layer. FS-DETR [7], the first few-shot detection transformer, extends DETR [9] to process arbitrary numbers of novel objects concurrently while supporting various examples from each class. Other approaches such as [27] prioritize salient representations through angle distance and magnitude dispersion across few-shot samples,  [93] leverages variational autoencoders for data augmentation, and  [61] learns geometry-aware features with an offline simplex equiangular tight frame classifier. Additionally,  [60] generates specific and representative prototypes for each query image, considering both support and query image features. Our method is a FSOD method based on region proposal generation and matching. Compared to previous FSOD methods, it is simple yet effective.
@@
Instance Detection. Instance detection involves identifying an unseen instance in a test image using corresponding templates. Some methods, e.g.,  [64, 2, 57], rely on pure 2D representations and matching techniques. Specifically, DTOID [64] employs global object attention and a local pose-specific branch for template-guided heatmap prediction. However, these methods may struggle with variations in 2D appearance caused by occlusion or pose changes. In contrast, VoxDet [49] utilizes explicit 3D knowledge from multi-view templates, providing geometry-invariant representations. We generate 2D robust instance embeddings from templates with DINOv2 [68, 20].
@@
Adapters for Pre-trained Models. The application of adapters atop large pre-trained models has emerged as a prominent strategy, yielding significant improvements across various tasks. Previous works have leveraged adapters to enhance few-shot image classification tasks [29, 99, 70, 101, 83]. Other works have demonstrated effectiveness in language model fine-tuning [37, 39, 100, 30]. Additional works exhibit the promise of adapters in vision tasks including detection and segmentation [17, 16, 71, 92, 23]. For novel instance detection and segmentation, we propose the Weight Adapter to enhance performance. The adapter assigns weights to the embeddings from frozen backbones. SENets [38] employ the Squeeze-and-Excitation (SE) blocks to distribute weights. In contrast, our adapter is an independent model with a structural difference from the SE block. For instance, the SE block lacks a ReLU layer prior to the sigmoid function.
3 Method
@@
Novel Instance Detection and Segmentation (NIDS), also referred to as open-world detection and segmentation, is a task to locate and label unseen instances within a query image. We will assume that each of the $N$ target instances is represented by $K$ template images $I_{T}\in\mathbb{R}^{K\times 3\times W\times H}$ and their corresponding segmentation masks. During inference, the output of each query image $I_{Q}\in\mathbb{R}^{3\times W\times H}$ provides the bounding boxes for detection and instance masks for segmentation of these instances.
3.1 Instance Embedding Generation Stage
@@
3.2 Object Proposal Stage
@@
3.3 Embedding Refinement via an Adapter
Most instance embeddings are well separated. However, some embeddings from distinct instances may cluster together. For example, background object embeddings might be similar to targets. Some target instances may resemble each other. To address this issue, we employ learnable adapters to refine the embeddings. We train these adapters using the InfoNCE loss [67, 15], aiming to bring the embeddings of the same instance closer together while separating the embeddings of different instances. This training only uses the template images.
@@ 3.3 Embedding Refinement via an Adapter
Given an instance embedding $f$ and an $MLP$ (Multi-Layer Perceptron), the CLIP-Adapter modifies the embedding according to the formula $f_{c}=\alpha\times MLP(f)+(1-\alpha)\times f$. In this equation, $f_{c}$ represents the adapted embedding, which is a linear combination of the transformed embedding $MLP(f)$ and the original embedding $f$. The residual ratio $\alpha$ is set as 0.6.
@@
Our Weight Adapter (WA). The CLIP-Adapter refines instance embeddings by adding a new feature vector. However, this new vector might not align with the feature space of original instance embeddings, which can lead to overfitting. Since only several hundred template embeddings of target instances are trained and refined in some datasets, the embeddings of background objects can be spoiled as the process disrupts the distances between target and background instance embeddings. Assuming the feature space of original embeddings is robust and effective, it is crucial to fine-tune the instance embeddings within this space by constraining the scope of the adapted embeddings. This refinement can be accomplished by simply applying weights to the original instance embeddings.
Hence, we introduce our Weight Adapter, a compact MLP network depicted in Fig. 3. The operation of this adapter is described through the following equations:
$w$ $=\text{sigmoid}(MLP(\beta f)),$ (1)
$f_{w}$ $=w\odot(\beta f).$ (2)
Here, $\beta$ is a positive scalar. We set $\beta$ as 10. $w$ is the weights computed by passing the scaled embedding $\beta f$ through the MLP and then applying a $sigmoid$ function. The output $f_{w}$ is the element-wise product of $w$ and the scaled embedding $\beta f$, resulting in the weighted embedding.
We include a ReLU layer before the $sigmoid$ function, ensuring that the sigmoid’s output is constrained to the range $[0.5,1)$. Subsequently, each refined embedding is close to its original counterpart. Additionally, $\beta$ allows for broader adjustments within the feature space. Given that cosine similarity is utilized to compare embeddings in the subsequent phase, our adapter alters the interrelationships among the values within each embedding. Further adapter details are presented in Fig. 3.
Figure 3: Comparison of two adapters. (left) CLIP-Adapter. (right) Our weight adapter.
3.4 Matching Stage
This stage provides each proposal $i$ with an instance ID $o_{i}$ as its label and a confidence score $s_{i}$.
@@ 3.4 Matching Stage
Initially, $Q$ proposal embeddings $E_{P}$ are matched with $N\times K$ template embeddings $E_{T}$ using cosine similarity. It yields a matrix of template scores with dimensions $Q\times N\times K$, as illustrated in Fig. 2. For each instance, we aggregate all $K$ template scores to derive a matrix of instance scores with dimensions $Q\times N$. We employ $Max$ as the aggregation function for optimal results.
@@
After matching, we acquire labeled proposals. Each is defined by $\{b_{i},M_{i},o_{i},s_{i}\}$. Here, $b_{i}$ is the bounding box of an instance $M_{i}$ is the modal mask, which covers the visible object surface [84], $o_{i}$ is the instance ID, and $s_{i}$ is the confidence score. A confidence score threshold $\delta$ can be used to remove incorrect proposals.
Table 1: Detection results on the high-resolution dataset. avg indicates that we are evaluating all images, including those from the easy and hard scenes. We train these two adapters with 2400 template embeddings.

### v3

@@ Abstract
Novel Instance Detection and Segmentation (NIDS) aims at detecting and segmenting novel object instances given a few examples of each instance. We propose a unified, simple, yet effective framework (NIDS-Net) comprising object proposal generation, embedding creation for both instance templates and proposal regions, and embedding matching for instance label assignment. Leveraging recent advancements in large vision methods, we utilize Grounding DINO and Segment Anything Model (SAM) to obtain object proposals with accurate bounding boxes and masks. Central to our approach is the generation of high-quality instance embeddings. We utilized foreground feature averages of patch embeddings from the DINOv2 ViT backbone, followed by refinement through a weight adapter mechanism that we introduce.
We show experimentally that our weight adapter can adjust the embeddings locally within their feature space and effectively limit overfitting in the few-shot setting. Furthermore, the weight adapter optimizes weights to enhance the distinctiveness of instance embeddings during similarity computation. This methodology enables a straightforward matching strategy that results in significant performance gains. Our framework surpasses current state-of-the-art methods, demonstrating notable improvements in four detection datasets. In the segmentation tasks on seven core datasets of the BOP challenge, our method outperforms the leading published RGB methods and remains competitive with the best RGB-D method. We have also verified our method using real-world images from a Fetch robot and a RealSense camera.  [^1] Project page with video, code and appendix: https://irvlutd.github.io/NIDSNet/
I Introduction
Novel Instance Detection and Segmentation (NIDS) is a crucial task in robot perception aimed at identifying and locating unseen instances in images or videos, given a few examples of each instance. Suppose that a robot needs to grasp a specific, novel object instance from a cluttered desk given only a small number of multi-view template images of the object. NIDS can provide the precise bounding box and mask of the target given a query image.
The current paradigm for NIDS typically encompasses the following steps: (1) generating proposals from a query image, (2) obtaining embeddings of the proposals and the object templates, and (3) matching the embeddings of the proposals with those of the templates for identification. Recent work [1, 2, 3, 4] has utilized various open-world detectors, e.g., the Segment Anything Model (SAM) [5] or FastSAM [6], to obtain object proposals. However, the use of open-world detectors often results in the generation of region-based proposals rather than actual object proposals. This misidentification of regions as objects can lead to object identification issues. For example, a single object may be divided into multiple proposals, or some background regions may be misclassified as foreground objects. These false alarms cause issues for the following identification stage.
To embed proposals and templates, some existing works, e.g., [2, 3, 4], adopt the $cls$ token of DINOv2 [7, 8]. [1] used a 3D voxel representation. Ideally, for each specific unseen instance, the template embeddings from different views should be similar to each other, but markedly different from the embeddings of other instances. However, in these previous works, the embeddings of different instances may remain similar.
@@ I Introduction
To address the limitations of existing methods, we propose a framework for NIDS, as depicted in Fig. 1. Initially, we utilize Grounding DINO [9] with the text prompt “objects” on a cluttered query image to obtain bounding boxes of foreground objects. This method capitalizes on the inherent objectness of Grounding DINO using specialized prompts. Subsequently, we employ SAM [5] to create masks within these bounding boxes, generating precise object proposals comprising both bounding boxes and masks. For instance embeddings, we first extract average foreground features [10] from the patch embeddings of the DINOv2 ViT backbone. We then apply adapters to enhance instance embeddings by clustering similar instances and distancing different ones.
@@
The adapters are trained using the template images only, where the InfoNCE loss [11, 12] is applied to the feature embeddings after the adapters. To refine these embeddings, the CLIP-Adapter [13] introduces a residual vector added to the original embedding, which risks overfitting with a few training examples. This addition can spoil and destabilize the embeddings of non-target objects, disrupting the entire feature space and causing the framework to misclassify non-targets as target instances. To mitigate this issue, we instead introduce a novel weight adapter (WA) that modifies the original embeddings by applying learned weights. Since the matching process relies on the standard cosine similarity, all embedding dimensions are treated equally, which may not sufficiently distinguish instances. The weights learned from our weight adapter emphasize the most relevant embedding channels. This produces more distinctive instance representations and improves performance by enhancing the discriminative capacity of embeddings. Finally, after applying the weight adapter to the embeddings, we employ a straightforward matching approach [14], e.g., using stable matching or $argmax$, to assign instance labels to these proposals from the query image.
@@
Our approach has been validated on four detection datasets, seven segmentation datasets, and in the real world for NIDS. The framework significantly surpasses existing state-of-the-art methods. It demonstrates substantial increases in average precision (AP), with gains of 22.3, 46.2, 10.3, and 24.0 across these detection datasets. Moreover, our method outperforms the best published RGB and RGB-D methods on instance segmentation tasks over the seven core datasets of the BOP challenge [15].
@@
  • We propose NIDS-Net, a unified framework for novel instance detection and segmentation that adapts pre-trained vision models for this task.
@@
  • We utilize the objectness of the Grounding DINO model and region recognition of the Segmentation Anything Model (SAM) to obtain object proposals.
@@
  • We introduce the Weight Adapter, a tool designed to refine embeddings within their feature space while preventing overfitting. This approach yields significant improvements as the adaptation of the feature space becomes more robust and stable.
II Related Work
@@
Pretrained Models. Large-scale pretraining has demonstrated utility across various downstream tasks. Works such as [16, 17, 7] emphasize large-scale image representation learning. DINOv2 [7] offers robust features to represent unseen instances. These foundational models offer diverse capabilities, with the challenge lying in effectively leveraging their wealth of knowledge for specific use cases, such as novel instance detection. In our work, we leverage pre-trained vision models for NIDS.
@@
Instance Detection. Instance detection identifies an unseen instance in a test image using corresponding templates. Some methods such as [18, 19, 20], rely on pure 2D representations and matching techniques. However, these methods may struggle with variations in 2D appearance caused by occlusion or pose changes. In contrast, VoxDet [1] utilizes explicit 3D knowledge from multi-view templates, providing geometry-invariant representations. We generate 2D robust instance embeddings from templates with DINOv2 [7].
@@
Adapters for Pre-trained Models. The application of adapters atop large pre-trained models has emerged as a prominent strategy, yielding significant improvements across various tasks. Previous works have leveraged adapters to enhance few-shot image classification tasks [13, 21, 22]. For novel instance detection and segmentation, we propose the Weight Adapter to enhance performance. The adapter assigns weights to the embeddings from frozen backbones. SENets [23] employ the Squeeze-and-Excitation (SE) blocks to distribute weights. In contrast, our adapter is an independent model with a structural difference from the SE block. For instance, the SE block lacks a ReLU layer prior to the sigmoid function.
III Method
@@
The NIDS task is to locate and label novel object instances within a query image, given a set of template images of these objects. We assume that each of the $N$ target instances is represented by $K$ template images $I_{T}\in\mathbb{R}^{K\times 3\times W\times H}$ and their corresponding segmentation masks. During inference, the output of each query image $I_{Q}\in\mathbb{R}^{3\times W\times H}$ provides the bounding boxes for detection and instance masks for segmentation of these instances.
III-A Instance Embedding Generation Stage
@@
III-B Object Proposal Stage
@@
III-C Embedding Refinement via an Adapter
Most instance embeddings are well separated. However, some embeddings from distinct instances may cluster together. For example, non-target object embeddings might be similar to targets. Some target instances may resemble each other. To address this issue, we employ learnable adapters to refine the embeddings. We train these adapters using the InfoNCE loss [11, 12], aiming to bring the embeddings of the same instance closer together while separating the embeddings of different instances. This training only uses the few-shot template images.
@@ III-C Embedding Refinement via an Adapter
As illusrated in Fig. 3 (Left), given an instance embedding f and an MLP (Multi-Layer Perceptron), the CLIP-Adapter modifies the embedding according to the formula $\mathbf{f_{c}}=\alpha\times\text{MLP}(\mathbf{f})+(1-\alpha)\times\mathbf{f}$. In this equation, $\mathbf{f_{c}}$ represents the adapted embedding, which is a linear combination of the transformed embedding $\text{MLP}(\mathbf{f})$ and the original embedding $\mathbf{f}$. The residual ratio $\alpha$ is set as 0.6.
Figure 3: (Left) CLIP-Adapter [13] (Right) Our introduced weight adapter
@@
Our Weight Adapter (WA). CLIP-Adapter refines instance embeddings by adding a new feature vector. However, this vector may not align with the original feature space, potentially causing overfitting. In datasets where only a few hundred template embeddings of target instances are refined, this adaptation can spoil non-target object embeddings by altering the relative distances between target and non-target instances. It leads the framework to misclassify non-targets as targets. Given the robustness and effectiveness of the original embedding space, it is essential to fine-tune instance embeddings within this space by constraining the adaptation. This can be achieved by applying weights to the original embeddings.
We propose the Weight Adapter, a compact MLP-based network structure illustrated in Fig. 3 (Right). This adapter operates according to the following equations:
$\mathbf{w}$ $=\text{sigmoid}(\text{MLP}(\beta\mathbf{f})),$ (1)
$\mathbf{f_{w}}$ $=\mathbf{w}\odot(\beta\mathbf{f}),$ (2)
where $\beta$ is a predefined positive scalar. $\mathbf{w}$ represents the weights derived by passing the scaled embedding $\beta\mathbf{f}$ through an MLP, followed by a sigmoid activation function. The resultant $\mathbf{f_{w}}$ is the element-wise multiplication of $\mathbf{w}$ and $\beta\mathbf{f}$, yielding the weighted embedding. An ReLU activation precedes the sigmoid in our architecture, restricting the output to the range [0.5, 1), which ensures that each adapted embedding retains proximity to its original value.
We will use the cosine similarity between two embeddings to measure their similarity. For two embeddings $\mathbf{q}$ and $\mathbf{k}$, their cosine similarity is defined as:
$\cos(\mathbf{q},\mathbf{k})=\frac{\mathbf{q}\cdot\mathbf{k}}{\|\mathbf{q}\|\;\|\mathbf{k}\|}.$ (3)
With our weight adapter, we modify the embeddings using respective weight vectors $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ as follows:
$\mathbf{q}^{\prime}=\mathbf{w}_{1}\odot(\beta\mathbf{q}),$ (4)
$\quad\mathbf{k}^{\prime}=\mathbf{w}_{2}\odot(\beta\mathbf{k}).$ (5)
Consequently, the cosine similarity between $\mathbf{q}^{\prime}$ and $\mathbf{k}^{\prime}$ is expressed as:
$\cos(\mathbf{q}^{\prime},\mathbf{k}^{\prime})=\frac{\sum_{i}w_{1,i}\,w_{2,i}\,q_{i}\,k_{i}}{\sqrt{\sum_{i}w_{1,i}^{2}q_{i}^{2}}\;\sqrt{\sum_{i}w_{2,i}^{2}k_{i}^{2}}}.$ (6)
Here, $\beta$ acts as a scaling factor since the embedding values can be small. We simply set $\beta$ to 10. It facilitates feature scaling, stabilizing gradients and enhancing convergence. The computed weights from the adapter, $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, ensure that embeddings of the same instance are closely aligned. These weights allow us to find the important dimensions for similarity computation. Moreover, our adapter is versatile and flexible because it can be integrated with any image encoder or embedding generation mechanism.
III-D Matching Stage
@@
This stage provides each proposal $i$ with an instance ID $o_{i}$ as its label and a confidence score $s_{i}$. Initially, $Q$ proposal embeddings $E_{P}$ are matched with $N\times K$ template embeddings $E_{T}$ using cosine similarity. It yields a matrix of template scores with dimensions $Q\times N\times K$, as illustrated in Fig. 2. For each instance, we aggregate all $K$ template scores to derive a matrix of instance scores with dimensions $Q\times N$. We employ $Max$ as the aggregation function for optimal results.
@@
After matching, we acquire labeled proposals. Each is defined by $\{b_{i},M_{i},o_{i},s_{i}\}$. Here, $b_{i}$ is the bounding box of an instance. $M_{i}$ is the modal mask, which covers the visible object surface [15]. $o_{i}$ is the instance ID, and $s_{i}$ is the confidence score. A confidence score threshold $\delta$ can be used to remove incorrect proposals.
TABLE I: Detection results on the high-resolution dataset. avg indicates that we are evaluating all images, including those from the easy and hard scenes. We train these two adapters with 2400 template embeddings.